# From Pixels to Torques with Linear Feedback

### v2

@@ Abstract
We demonstrate the effectiveness of simple observer-based linear feedback policies for “pixels-to-torques” control of robotic systems using only a robot-facing camera. Specifically, we show that the matrices of an image-based Luenberger observer (linear state estimator) for a “student” output-feedback policy can be learned from demonstration data provided by a “teacher” state-feedback policy via simple linear-least-squares regression. The resulting linear output-feedback controller maps directly from high-dimensional raw images to torques while being amenable to the rich set of analytical tools from linear systems theory, allowing us to enforce closed-loop stability constraints in the learning problem. We also investigate a nonlinear extension of the method via the Koopman embedding. Finally, we demonstrate the surprising effectiveness of linear pixels-to-torques policies on a cartpole system, both in simulation and on real-world hardware. The policy successfully executes both stabilizing and swing-up trajectory tracking tasks using only camera feedback while subject to model mismatch, process and sensor noise, perturbations, and occlusions.
@@ 1 Introduction
Both model-based [48, 47, 34, 25, 51, 32, 43, 27, 45] and learning-based [38, 36, 37, 29, 40, 46, 1] policies have demonstrated impressive control results in robotics using vision-based sensory feedback. However, both suffer from inherent drawbacks: neural-network policies trained from images can require large amounts of data, which often must be generated in simulation before being transferred to a real robot. Meanwhile, model-based policies require separate machine-vision algorithms to extract features for downstream state estimation or model learning. Recently, data-driven switched linear control policies have been applied directly to image feedback with promising results [41]. However, this was limited to a quasi-static setting. We aim to apply data-driven linear feedback policies to the control of dynamic robotic systems directly from images as shown in Fig. 1.
@@
We demonstrate that it is possible to perform “pixels-to-torques” control of robotic systems with simple observer-based linear feedback policies in both simulation and on real hardware. Specifically, we show that an image-based linear output-feedback policy can be designed by learning a Luenberger observer (i.e., a linear state estimator) via linear-least-squares (LLS) regression over supervised demonstration data. The result is a pixels-to-torques policy that is interpretable and can be analyzed using the rich set of tools from classical linear systems theory, all while inferring states directly from high-dimensional images. We leverage this interpretability to promote stability of the closed-loop policy via a convex extension of the LLS problem. The resulting linear output-feedback policy can achieve good performance while learning from small amounts of data, which we demonstrate on a classic cartpole system. Specifically, the policy is able to avoid sim-to-real transfer by learning directly from hardware data and successfully stabilizes the system while being robust to perturbations, unmodeled dynamics, and even visual occlusions. To summarize, our contributions are:
  • A data-efficient “student-teacher” methodology for learning linear visual output-feedback policies.
@@
  • Demonstration of linear “pixels-to-torques” control on a real-world cartpole system.
The remainder of the paper is organized as follows: In Section 2, we provide a broad survey of related works on image-based control and estimation in robotics. In Section 3, we provide an overview of topics from linear control, state estimation, and Koopman-operator theory. In Section 4, we describe our linear-policy framework as well as the student-teacher-based method developed to learn a Luenberger observer. We also introduce a Koopman-based extension for nonlinear control settings. Section 5 presents experimental results demonstrating the method’s performance on a cartpole system both in simulation and on real-world hardware. Lastly, we summarize our conclusions in Section 6.
@@ 2.2 Learning Policies with Image Feedback
Learning control policies directly from visual inputs has been a widely used practice in the deep-RL community  [36, 37, 1, 29]. While initially performed within virtual environments [36, 37], deep-neural-network image-feedback policies have quickly been extended to real-world robotics tasks, including manipulation via diffusion policies [8] and quadruped locomotion over various terrains with egocentric vision [1]. These methods, while demonstrating impressive results, are limited by the need for large amounts of training data. To scale to learning on real-world hardware, Levine et. al. [29] proposed a guided policy search method for efficiently learning an end-to-end pixels-to-torques policy for manipulation tasks. Specifically, “guiding” linear-Gaussian controllers with privileged full-state feedback are optimized and used to provide supervised training data for learning the pixels-to-torques policy. Rather than learning a deep end-to-end policy, our method learns a simple Luenberger observer from the guided supervised data to achieve image-to-state estimation for a downstream state-feedback controller.
@@ 2.3 Vision-Based Pose and State Estimation
Deep learning has been widely applied to vision-based pose and state estimation across many robotics applications [48, 39, 28, 50, 52, 26]. Applications include both external feature detection for a downstream controller to track [39, 48] and estimation of the robot’s own pose or state [52, 26, 28, 50]. A commonality across these learning-based methods is the need for large amounts of data, necessitating data synthesis and sim-to-real transfer. As a result, Lu et. al. looked to differentiable rendering for both robotic pose and state estimation from images [31, 30]. Specifically, Lu et. al. rendered predicted images of the robot and calculated the corresponding Jacobians for an extended Kalman filter, resulting in online robot-state estimation from a single camera [31]. However, this method requires a differentiable rendering model of the robot to generate predicted observations. Therefore, we aim to determine if it’s possible to directly learn an image-based Luenberger observer instead of relying on an explicit observation model.
@@ 2.4 Effectiveness of Linearization
Over many decades, linear controllers and state estimators have proven to be powerful tools within the controls community thanks to their simplicity, ease of analysis, and surprising robustness [23, 22]. Recently, under the label of state-space models (SSMs), linear models have also become popular in the deep-learning community [17, 19, 18]. Specifically, Gu et. al. [17, 19] demonstrated that structured linear models in the latent spaces of deep networks can model long-range dependencies found in many applications like speech recognition, language processing, and genomics [18]. The most recent implementations have produced performance matching or exceeding that of widely-adopted transformers while being much more computationally efficient [17]. We note the strong similarity of SSMs to deep-Koopman methods [33], specifically in their learning of linear dynamics in network latent spaces.
In addition to deep-learning methods, non-neural linear alternatives have also been shown to be surprisingly effective in robotics [21, 5, 31] and vision-based applications [41, 32]. By applying approximate Koopman operators [24], both Haggerty et. al. [21] and Bruder et. al. [5] were able to model highly nonlinear, high-dimensional soft manipulators with no neural components. Within the context of quasi-static pile manipulation, Suh et. al. [41] showed that switched linear visual-foresight models can outperform deep counterparts in both prediction error and closed-loop control performance. However, applying these model-learning methods directly to image-based observations of open-loop unstable dynamical systems can be difficult. Koopman liftings of high-dimensional images is computationally expensive, and linear visual-foresight models has been restricted to quasi-static systems. Therefore, we aim to determine if it is possible to avoid the development of a Koopman dynamics model by directly learning an observer-based linear feedback policy.
@@ 3.1 Linear Control
Using the dynamics described by (1), we can design a linear state-feedback controller to drive the robot to the origin:
$u_{k}=-K\hat{x}_{k},$ (2)
where $K\in\mathbb{R}^{m\times n}$ is the controller gain matrix. As a particular example, K could be determined by solving an LQR problem using the Ricatti equation [22].
@@
where we can see that $x_{k}$ will converge to the origin if $A-BK$ is stable. This stability condition for the closed-loop dynamics can be expressed as,
@@ 3.2 Linear State Estimation
An observer can then be designed to first predict states using the dynamics, (1), before correcting the predictions using sensor observations to provide the final full-state estimates. This is the core idea behind a Luenberger observer (i.e. linear state estimator):
@@
where $\hat{x}_{k+1}$ are the estimated state output by the observer at time step $k+1$; $\hat{y}_{k+1}$ are the corresponding predicted observations; $y_{k+1}$ are the corresponding real-world observations (i.e., sensor readings on hardware); and $L\in\mathbb{R}^{n\times l}$ is the observer gain. As an example, a Kalman filter can be used to design $L$ to properly estimate $\hat{x}_{k+1}$ under Gaussian noise. We refer to the combined controller-and-observer system as an output-feedback policy.
Similarly to $K$, $L$ can also be calculated to drive the state estimation errors, $x_{k}-\hat{x}_{k}$, asymptotically to zero for a given linear feedback controller. We first account for the controller by substituting (2) into (6):
$\hat{x}_{k+1}=(A-BK)\hat{x}_{k}+L(y_{k+1}-\hat{y}_{k+1}).$ (7)
Then, we substitute (1) and (5) into (7) to further simplify the Luenberger observer,
@@
This allows us to analyze the stability of the Luenberger observer in a similar manner to the linear state-feedback controller as described by (4). Specifically, we can express the condition for asymptotically-converging state estimation errors as:
@@
We will define the matrix $A^{LK}$ as:
@@ 3.3 Koopman-Operator Theory
where $z\in\mathbb{R}^{N_{z}}$. Likewise, $\phi$ is constructed in such a way that the “unlifting” is linear:
@@
Figure 2: Overview of learning a linear output-feedback policy for successful pixels-to-torques control of a robotic system. In this work, a linear-quadratic-Gaussian (LQG) “teacher” policy is designed to control a cartpole system. Teacher-demonstration data is collected as trajectories of the robot’s estimated states, $\hat{X}_{1:N}$; control inputs, $U_{1:N-1}$; and corresponding images from a robot-facing camera, $Y^{p}_{2:N}$. Subsequent linear-least-squares (LLS) regression is performed over the data to determine the parameters, $\theta$, of a “student” policy’s pixel-based Luenberger observer (i.e., linear state estimator). The student policy’s controller can also be learned separately or be the same as the teacher’s, as is the case in this work. The solid and dotted lines indicate processes that were performed online and offline respectively.
@@ 4 Methodology
For a robot controlled via image-based feedback, the development of an appropriate observation model, (5), requires a differentiable renderer [30] that we may not have access to. In addition, performing a gain-calculation procedure, such as solving the Ricatti equation, may be too computationally expensive to perform using dynamics models in pixel space [41]. Therefore, we directly learn the parameters of a linear feedback policy’s Luenberger observer, which uses images from a robot-facing camera as sensory feedback. An overview of our method is shown in Fig. 2.
@@ 4.1 Pixel-Based Luenberger Observer
We begin by restating the Luenberger observer. Specifically, we substitute (1) and (5) into (6) to express the Luenberger observer in terms of just the predicted robot states, $\hat{x}$; robot control inputs, $u$; and output observations in the form of high-dimensional, raw pixel values from a robot-facing camera, $y^{p}$:
@@ 4.2 Learning a Pixel-Based Linear Output-Feedback Policy
$W=\begin{bmatrix}\hat{X}_{1:N}\\ U_{1:N-1}\\ Y^{p}_{2:N}\\ 1\end{bmatrix}=\begin{bmatrix}\hat{x}_{1}&\hat{x}_{2}&\dots&\hat{x}_{N-1}\\ u_{1}&u_{2}&\dots&u_{N-1}\\ y^{p}_{2}&y^{p}_{3}&\dots&y^{p}_{N}\\ 1&1&\dots&1\end{bmatrix},\quad\hat{X}_{2:N}=\begin{bmatrix}\hat{x}_{2}&\hat{x}_{3}&\dots&\hat{x}_{N}\end{bmatrix}.$ (18)
@@
We additionally introduce sparsity-promoting $L_{1}$ regularization to reduce overfitting,
@@
By combining the teacher’s state-feedback controller with the newly designed Luenberger observer, the student output-feedback policy can then perform closed-loop pixels-to-torques control as shown in Fig. 2. We note that regression can be separately performed over states and control histories to also learn a gain, $K$, for a new student controller. However, if the teacher is a linear feedback policy, this is equivalent to using the teacher’s controller.
@@ 4.3 Enforcing Stability with a Linear-Matrix-Inequality
By leveraging linear systems theory, we introduce a convex constraint to the observer-learning problem to ensure that state-estimation errors are guaranteed not to diverge. Specifically, we note that the stability condition described by (9) is equivalent to the following spectral-norm condition,
@@
$W=\begin{bmatrix}\hat{X}_{1:N}\\ Y^{p}_{2:N}\\ 1\end{bmatrix},\quad\theta=\begin{bmatrix}A^{LK}&L&d\end{bmatrix}.$
@@ 4.4 Extension to Nonlinear Systems via Koopman Embedding
and $\theta$ represents the concatenated matrices of the Koopman Luenberger:
@@ 5.1 Simulation Setup
Figure 3: Stabilization performance vs. training trajectories of a pixel-based linear output-feedback policy tasked with stabilizing a cartpole using a robot-facing camera. Test stabilizations from 50 different initial conditions were evaluated with stabilization error defined as the $L^{2}$ error of the final state w.r.t the upright goal state. The median error is shown as a thick line, while the shaded regions represent the 5% to 95% bounds. The corresponding success rate of stabilization from the 50 initial conditions is also shown.

### v3

@@ Abstract
We demonstrate the effectiveness of simple observer-based linear feedback policies for “pixels-to-torques” control of robotic systems using only a robot-facing camera. Specifically, we show that the matrices of an image-based Luenberger observer (linear state estimator) for a “student” output-feedback policy can be learned from demonstration data provided by a “teacher” state-feedback policy via simple linear-least-squares regression. The resulting linear output-feedback controller maps directly from high-dimensional raw images to torques while being amenable to the rich set of analytical tools from linear systems theory, allowing us to enforce closed-loop stability constraints in the learning problem. We also investigate a nonlinear extension of the method via the Koopman embedding. Finally, we demonstrate the surprising effectiveness of linear pixels-to-torques policies on a cartpole system, both in simulation and on real hardware. The policy successfully executes both stabilizing and swing-up trajectory-tracking tasks using only camera feedback while subject to model mismatch, process and sensor noise, perturbations, and occlusions. Open-source code for all experiments can be found here: https://roboticexplorationlab.org/projects/linear_pixels_to_torques.html
@@ 1 Introduction
Both model-based [48, 47, 34, 25, 51, 32, 43, 27, 45] and learning-based [38, 36, 37, 29, 40, 46, 1] control policies have demonstrated impressive results in robotics using vision-based sensory feedback. However, both suffer from inherent drawbacks: neural-network policies trained from images can require large amounts of data, which often must be generated in simulation before being transferred to a real robot. Meanwhile, model-based policies require separate machine-vision algorithms to extract features for downstream state estimation or model learning. Recently, data-driven switched linear control policies have been applied directly to image feedback with promising results [41]. However, this was limited to a quasi-static setting. We aim to apply data-driven linear feedback policies to the control of dynamic robotic systems directly from images as shown in Fig. 1.
@@
We demonstrate that it is possible to perform “pixels-to-torques” control of robotic systems with simple observer-based linear feedback policies in both simulation and on real hardware. Specifically, we show that an image-based linear output-feedback policy can be designed by learning a Luenberger observer (linear state estimator) via linear-least-squares (LLS) regression over supervised demonstration data. The result is a pixels-to-torques policy that is interpretable and can be analyzed using the rich set of tools from classical linear systems theory, all while inferring states directly from high-dimensional images. We leverage this interpretability to promote stability of the closed-loop policy via a convex extension of the LLS problem. The resulting linear output-feedback policy can achieve good performance while learning from small amounts of data, which we demonstrate on a classic cartpole system. Specifically, the policy is able to avoid sim-to-real transfer issues by learning directly from hardware data, and it successfully stabilizes the system while being robust to perturbations, unmodeled dynamics, and visual occlusions. To summarize, our contributions are:
  • A data-efficient “student-teacher” methodology for learning linear, visual output-feedback policies.
@@
  • Demonstration of linear “pixels-to-torques” control on a real-world cartpole.
The remainder of the paper is organized as follows: In Section 2, we provide a broad survey of related works on image-based control and estimation in robotics. In Section 3, we provide an overview of topics from linear control, state estimation, and Koopman-operator theory. In Section 4, we describe our linear policy framework as well as the student-teacher method developed to learn a Luenberger observer. We also introduce a Koopman-based extension for nonlinear control settings. Section 5 presents experimental results demonstrating the method’s performance on a cartpole system both in simulation and on real-world hardware. Lastly, we summarize our conclusions in Section 6.
@@ 2.2 Learning Policies with Image Feedback
Learning control policies directly from visual inputs has been a widely used practice in the deep-RL community  [36, 37, 1, 29]. While initially performed within virtual environments [36, 37], deep-neural-network image-feedback policies have quickly been extended to real-world robotics tasks, including manipulation via diffusion policies [8] and quadruped locomotion over various terrains with egocentric vision [1]. These methods, while demonstrating impressive results, are limited by the need for large amounts of training data. To scale to learning on real-world hardware, Levine et. al. [29] proposed a guided policy search method for efficiently learning an end-to-end pixels-to-torques policy for manipulation tasks. Specifically, “guiding” linear controllers with privileged full-state feedback are optimized and used to provide supervised training data for learning the pixels-to-torques policy. Rather than learning a deep end-to-end policy, our method learns a simple Luenberger observer from the guided supervised data to achieve image-to-state estimation for a downstream state-feedback controller.
@@ 2.3 Vision-Based Pose and State Estimation
Deep learning has been widely applied to vision-based pose and state estimation across many robotics applications [48, 39, 28, 50, 52, 26]. Applications include both external feature detection for a downstream controller to track [39, 48] and estimation of the robot’s own pose or state [52, 26, 28, 50]. A commonality across these learning-based methods is the need for large amounts of data, necessitating data synthesis and sim-to-real transfer. As a result, Lu et. al. looked to differentiable rendering for both robot pose and state estimation from images [31, 30]. Specifically, Lu et. al. rendered predicted images of the robot and calculated the corresponding Jacobians for an extended Kalman filter, resulting in online robot-state estimation from a single camera [31]. However, this method requires a differentiable rendering model of the robot to generate predicted observations. Therefore, we aim to determine if it’s possible to directly learn an image-based Luenberger observer instead of relying on an explicit observation model.
@@ 2.4 Effectiveness of Linearization
Over many decades, linear controllers and state estimators have proven to be powerful tools within the controls community thanks to their simplicity, ease of analysis, and often surprisingly good performance on nonlinear systems [23, 22]. Recently, under the label of state-space models (SSMs), linear models have also become popular in the deep-learning community [17, 19, 18]. Specifically, Gu et. al. [17, 19] demonstrated that structured linear models in the latent spaces of deep networks can model long-range dependencies found in many applications like speech recognition, language processing, and genomics [18]. The most recent implementations have produced performance matching or exceeding that of widely-adopted transformers while being much more computationally efficient [17]. We note the strong similarity of SSMs to deep-Koopman methods [33], specifically in their learning of linear dynamics in a latent space.
In addition to deep-learning methods, non-neural linear alternatives have also been shown to be surprisingly effective in robotics [21, 5, 31] and vision-based applications [41, 32]. By applying approximate Koopman operators [24], both Haggerty et. al. [21] and Bruder et. al. [5] were able to model highly nonlinear, high-dimensional soft manipulators with no neural components. Within the context of quasi-static pile manipulation, Suh et. al. [41] showed that switched linear visual-foresight models can outperform deep counterparts in both prediction error and closed-loop control performance. However, applying these model-learning methods directly to image-based observations of open-loop unstable dynamical systems can be difficult. Koopman liftings of high-dimensional images are computationally expensive, and linear visual-foresight models have been restricted to quasi-static systems. Therefore, we aim to determine if it is possible to avoid the development of a Koopman dynamics model by directly learning an observer-based linear feedback policy.
@@ 3.1 Linear Control
Using the dynamics described by (1), we can design a linear state-feedback controller to drive the robot state to the origin:
$u_{k}=-Kx_{k},$ (2)
where $K\in\mathbb{R}^{m\times n}$ is the controller gain matrix. As a particular example, K could be determined by solving an LQR problem using the Riccati equation [22].
@@
where we can see that $x_{k}$ will converge to the origin if $A-BK$ is stable:
@@ 3.2 Linear State Estimation
An observer can then be designed to first predict states using the dynamics, (1), before correcting the predictions using sensor observations to provide the final full-state estimates. This is the core idea behind a Luenberger observer (linear state estimator):
@@
where $\hat{x}_{k+1}$ is the estimated state output by the observer at time step $k+1$; $\hat{y}_{k+1}$ are the corresponding predicted observations; $y_{k+1}$ are the corresponding real-world observations (i.e., sensor measurements); and $L\in\mathbb{R}^{n\times l}$ is the observer gain. $L$ is often designed with measurement noise and model uncertainty in mind, while the estimator’s internal state $\hat{x}_{k}$ maintains a memory of aggregated information from previous time steps [22]. As an example, a Kalman filter can be used to design $L$ to optimally estimate $\hat{x}_{k}$ when all uncertainties in the system correspond to additive Gaussian noise.
The state estimates, $\hat{x}_{k}$, from the Luenberger observer at time step $k$ can be subsequently passed to a linear state-feedback controller to calculate the corresponding control inputs,
$u_{k}=-K\hat{x}_{k}.$ (7)
We refer to this combined controller-observer system as an output-feedback policy with an internal state, $\hat{x}$.
Similarly to $K$, $L$ can also be calculated to drive the state estimation errors, $x_{k}-\hat{x}_{k}$, asymptotically to zero for a given linear feedback controller. We first express $\hat{y}_{k+1}$ in terms of $\hat{x}_{k}$ and $u_{k}$ by substituting the prediction term of (6) into (5):
$\hat{y}_{k+1}=C(A\hat{x}_{k}+Bu_{k}).$ (8)
Then, we substitute (7) and (8) into (6) to further simplify the observer,
@@
This allows us to analyze the stability of the Luenberger observer in a similar manner to the linear state-feedback controller as described by (4). Specifically, we can express the condition for asymptotically converging state-estimation errors as:
@@
For clarity, we will define the matrix $A^{LK}$:
@@ 3.3 Koopman-Operator Theory
where $z\in\mathbb{R}^{j}$. Likewise, $\phi$ is constructed in such a way that the “unlifting” is linear:
@@
Figure 2: Overview of learning a linear output-feedback policy for pixels-to-torques control. In this work, a linear-quadratic-Gaussian (LQG) “teacher” policy is designed to control a cartpole system. Teacher demonstration data is collected as trajectories of the robot’s estimated states, $\hat{X}_{1:N}$; control inputs, $U_{1:N-1}$; and corresponding images from a robot-facing camera, $Y^{p}_{2:N}$. Subsequent linear-least-squares (LLS) regression is performed over the data to determine the parameters, $\theta$, of a “student” policy’s pixel-based Luenberger observer (linear state estimator). The student policy’s controller can also be learned separately or cloned from the teacher’s, as is the case in this work. The solid and dotted lines indicate processes that were performed online and offline respectively.
@@ 4 Methodology
For a robot controlled via image-based feedback, the development of an appropriate observation model, (5), requires a differentiable renderer [30] that we may not have access to. In addition, performing a gain-calculation procedure, such as solving the Riccati equation, may be too computationally expensive to perform using dynamics and observation models in pixel space [41]. Therefore, we directly learn the parameters of a linear feedback policy’s Luenberger observer, which uses images from a robot-facing camera as sensory feedback. An overview of our method is shown in Fig. 2.
@@ 4.1 Pixel-Based Luenberger Observer
We begin by restating the Luenberger observer. Specifically, we substitute (8) into (6) to express the Luenberger observer in terms of just the predicted robot states, $\hat{x}$; robot control inputs, $u$; and output observations in the form of high-dimensional, raw pixel values from a robot-facing camera, $y^{p}$:
@@ 4.2 Learning a Pixel-Based Linear Output-Feedback Policy
$W=\begin{bmatrix}\hat{X}_{1:N-1}\\ U_{1:N-1}\\ Y^{p}_{2:N}\\ 1\end{bmatrix}=\begin{bmatrix}\hat{x}_{1}&\hat{x}_{2}&\dots&\hat{x}_{N-1}\\ u_{1}&u_{2}&\dots&u_{N-1}\\ y^{p}_{2}&y^{p}_{3}&\dots&y^{p}_{N}\\ 1&1&\dots&1\end{bmatrix},\quad\hat{X}_{2:N}=\begin{bmatrix}\hat{x}_{2}&\hat{x}_{3}&\dots&\hat{x}_{N}\end{bmatrix}.$ (19)
@@
We additionally introduce sparsity-promoting $L_{1}$ regularization to reduce overfitting:
@@
By combining the teacher’s state-feedback controller with the newly designed Luenberger observer, the student output-feedback policy can perform closed-loop pixels-to-torques control as shown in Fig. 2. We note that regression can be separately performed over states and control histories to also learn a gain, $K$, for a new student controller. However, if the teacher is a linear feedback policy, this is equivalent to using the teacher’s controller.
@@ 4.3 Enforcing Stability with a Linear-Matrix-Inequality
By leveraging existing techniques from linear systems theory [22, 4], we introduce a convex constraint to the observer-learning problem to ensure that state-estimation errors are guaranteed not to diverge. Specifically, we note that the stability condition described by (10) is equivalent to the following spectral-norm condition,
@@
$W=\begin{bmatrix}\hat{X}_{1:N-1}\\ Y^{p}_{2:N}\\ 1\end{bmatrix},\quad\theta=\begin{bmatrix}A^{LK}&L&d\end{bmatrix}.$
@@ 4.4 Extension to Nonlinear Systems via Koopman Embedding
and $\theta$ represents the concatenated matrices of the Koopman Luenberger observer:
@@
Figure 3: Stabilization performance vs. training trajectories of a pixel-based linear output-feedback policy tasked with stabilizing a cartpole using a robot-facing camera. Test stabilizations from 100 different initial conditions were evaluated with stabilization error defined as the $L^{2}$ error of the final state w.r.t the upright goal state. The median error is shown as a thick line, while the shaded regions represent the 5% to 95% bounds. The corresponding success rate of stabilization from the 100 initial conditions is also shown.
5 Experimental Results
5.1 Simulation Setup